# K-theory and localizing invariants of large categories

### v2

@@ Appendix E Phantom maps, pure injectivity and Adams representability
is zero. If moreover is a Serre subcategory of then one can show that any fp-injective object of has injective dimension at most
Now, for any dualizable category an almost -module is almost cohomological if and only if it is fp-injective in If is -compact, it can be shown that is a Serre subcategory in This gives an alternative (intrinsic) way to show that has injective dimension at most for any
@@
If is a Grothendieck abelian category with (AB6) and (AB4*), then one can define flat objects of intrinsically. Namely, an object is flat if for any compact morphism and for any object the map is zero.
If is a dualizable category, then a right almost -module is almost cohomological if and only if it is flat in
@@ Appendix F Analogy between dualizable categories and compact Hausdorff spaces
Denote by the category of compact Hausdorff spaces, and denote by the full subcategory of profinite sets. We have a functor By Corollary 6.15, the category is compactly generated if and only if is profinite.
It turns out that a lot of statements about dualizable categories discussed in this paper have very natural analogues for compact Hausdorff spaces. The table below summarizes this analogy. Here for a topological space we denote by its set of quasi-components, with the quotient topology. Recall that quasi-components of are equivalence classes of points, where if any open-closed subset of containing also contains
@@
Epimorphisms in are maps with a dense image. It follows that satisfies weak (AB4*), i.e. the class of epimorphisms is closed under products.
Filtered colimits in are obtained by taking the completion (of the separation of) of the naive filtered colimit. Here the norm on the naive (filtered) colimit is defined by
To show that (AB6) holds in recall the presentable compactly assembled category (of prenormed vector spaces and contractive maps) from Example 1.34. Since is compactly assembled, it satisfies (AB6). It remains to observe that the completion functor commutes with all colimits and with infinite products.∎
@@
The functor is conservative by Urysohn’s lemma. By Tietze extension theorem, the functor detects epimorphisms (and also detects regular epimorphisms).
Now, the product of in is given by – Stone-Čech compactification of the disjoint union. For any topological space continuous functions on are identified with bounded continuous functions on It follows that the functor commutes with products.
Finally, consider a codirected system where is directed, and put By Stone-Weierstrass theorem, the union of the images is a dense subalgebra. It remains to observe that the map from the naive colimit to is isometric. Indeed, this follows from the fact that for any and for any neighborhood of there is some such that It follows that the functor commutes with filtered colimits. ∎
@@
Consider a family of profinite sets We need to show that the space is totally disconnected. It suffices to show that the locally constant -valued functions on are dense in the space of all continuous functions.
Consider a function corresponding to a family of continuous functions such that for some we have Take some and choose an integer Since is profinite, we can find a locally constant function such that and takes values in the finite set Then the corresponding function is also locally constant and ∎
@@
To see that the object is -compact in it suffices to observe that -filtered colimits in can be computed naively, i.e. the naive colimit is already complete (and separated). By Urysohn’s lemma, the functor is conservative. Hence, the object generates the category by colimits.
Now, the full subcategory of metrizable spaces in contains the final object and is closed under fiber products and countable products. Hence, metrizable spaces in are closed under -small colimits. It remains to show that they are generated by via -small colimits.
Let be a metrizable compact Hausdorff space. Choose an injective map Denote by the pushout Then is also metrizable, and we choose an injective map Composing it with the two maps we obtain two self-maps The equalizer is isomorphic to ∎
@@
Consider a codirected system and put for Passing to Boolean algebras, we need to show that the map is an isomorphism. Given an open-closed consider its characteristic function Then for some we have a continuous function such that Then is an open-closed subset of and its preimage in equals
Suppose that is an open-closed subset such that Then hence for some we have i.e. ∎
@@
Consider the surjective map Put and Then
Put Then the diagram is the limit of but is a one-element set for Hence, the sequential limits do not commute with coequalizers in ∎
@@ Appendix G K-theory of products of exact infinity-categories
In this section we prove that -theory of exact -categories commutes with infinite products.
@@ G.1. Products of additive infinity-categories
Given a small additive -category we denote by its stabilization. By [Bon, Theorem 5.3.1, Proposition 6.2.1], if is Karoubi complete, then so is Given we denote by the full subcategory generated by We put and similarly
The following result is due to A. Córdova Fedeli.
@@
Since we will need an analogous statement for -structures, we recall the proof from [Cor23] with some simplifications.
We introduce some notation. Put Given functions we put
@@
2) Analogous statement holds for the same functors
@@
Then and the map becomes an isomorphism in This proves that is essentially surjective.
Let now Let be objects of and denote by and their images in
Fullness of Consider a map represented by a roof where and Take such that and Choose and the maps such that Then we can choose factorizations and automatically
Putting we have and Therefore, the roof represents a morphism such that
Faithfulness of Consider a map such that We may assume that can be lifted to a morphism Then there exists a map such that and Arguing as above, we can construct a factorization of such that and Hence ∎
@@ G.2. Products of stable categories with bounded -structures
We will need the following analogue of Theorem G.1. Let be a family of stable categories with bounded -structures. Denote by the full subcategory formed by objects such that the objects have uniformly bounded homological amplitude.
@@
The proof is completely analogous to the proof of Theorem G.1 with minor changes, which we indicate.
For a stable category with a bounded -structure and for integers we denote by the full subcategory formed by objects whose homology is concentrated in degrees between and and similarly for
Analogously to Lemma G.3, one shows that we have an isomorphism
Analogously to Lemma G.4, one proves that for any function the functor is an equivalence.
Then one concludes that proving the theorem.∎
@@ G.3. Products of exact -categories
[Bar15, Definition 3.1] An exact -category is a triple where is an additive -category and are subcategories such that the following conditions hold.
1) Every morphism (resp. ) is in (resp. ).
2) Pushouts of morphisms in exist, and the class of morphisms in is closed under pushouts. Pullbacks of morphisms in exist and the class of morphisms in is closed under pullbacks.
3) For a commutative square
the following are equivalent:
a) the square is coCartesian, and
b) the square is Cartesian, and
@@
where and A total cofiber of such a square is by definition a primitive acyclic object. The full subcategory of acyclic objects is the stable subcategory generated by the primitive acyclic objects. The stable hull of is the category we follow the notation of [Kle22].
@@
Note that if is a triangulated category with a -structure, and is a weak Serre subcategory of the heart, then the full triangulated subcategory generated by has a bounded -structure with the heart In particular, the functor is -exact.

### v3

@@ Appendix E Phantom maps, pure injectivity and Adams representability
is zero. If moreover is a Serre subcategory of then one can show that any fp-injective object of has injective dimension at most Now, for any dualizable category an almost -module is almost cohomological if and only if it is fp-injective in If is -compact, it can be shown that is a Serre subcategory in This gives an alternative (intrinsic) way to show that has injective dimension at most for any
@@
If is a Grothendieck abelian category with (AB6) and (AB4*), then one can define flat objects of intrinsically. Namely, an object is flat if for any compact morphism and for any object the map is zero. If is a dualizable category, then a right almost -module is almost cohomological if and only if it is flat in
@@ Appendix F Analogy between dualizable categories and compact Hausdorff spaces
Denote by the category of compact Hausdorff spaces, and denote by the full subcategory of profinite sets. We have a functor By Corollary 6.16, the category is compactly generated if and only if is profinite. It turns out that a lot of statements about dualizable categories discussed in this paper have very natural analogues for compact Hausdorff spaces. The table below summarizes this analogy. Here for a topological space we denote by its set of quasi-components, with the quotient topology. Recall that quasi-components of are equivalence classes of points, where if any open-closed subset of containing also contains
@@
Epimorphisms in are maps with a dense image. It follows that satisfies weak (AB4*), i.e. the class of epimorphisms is closed under products. Filtered colimits in are obtained by taking the completion (of the separation of) of the naive filtered colimit. Here the norm on the naive (filtered) colimit is defined by
To show that (AB6) holds in recall the presentable compactly assembled category (of seminormed vector spaces and contractive maps) from Example 1.33. Since is compactly assembled, it satisfies (AB6). It remains to observe that the completion functor commutes with all colimits and with infinite products.∎
@@
The functor is conservative by Urysohn’s lemma. By Tietze extension theorem, the functor detects epimorphisms (and also detects regular epimorphisms). Now, the product of in is given by – Stone-Čech compactification of the disjoint union. For any topological space continuous functions on are identified with bounded continuous functions on It follows that the functor commutes with products. Finally, consider a codirected system where is directed, and put By Stone-Weierstrass theorem, the union of the images is a dense subalgebra. It remains to observe that the map from the naive colimit to is isometric. Indeed, this follows from the fact that for any and for any neighborhood of there is some such that It follows that the functor commutes with filtered colimits. ∎
@@
Consider a family of profinite sets We need to show that the space is totally disconnected. It suffices to show that the locally constant -valued functions on are dense in the space of all continuous functions. Consider a function corresponding to a family of continuous functions such that for some we have Take some and choose an integer Since is profinite, we can find a locally constant function such that and takes values in the finite set Then the corresponding function is also locally constant and ∎
@@
To see that the object is -compact in it suffices to observe that -filtered colimits in can be computed naively, i.e. the naive colimit is already complete (and separated). By Urysohn’s lemma, the functor is conservative. Hence, the object generates the category by colimits. Now, the full subcategory of metrizable spaces in contains the final object and is closed under fiber products and countable products. Hence, metrizable spaces in are closed under -small colimits. It remains to show that they are generated by via -small colimits. Let be a metrizable compact Hausdorff space. Choose an injective map Denote by the pushout Then is also metrizable, and we choose an injective map Composing it with the two maps we obtain two self-maps The equalizer is isomorphic to ∎
@@
Consider a codirected system and put for Passing to Boolean algebras, we need to show that the map is an isomorphism. Given an open-closed consider its characteristic function Then for some we have a continuous function such that Then is an open-closed subset of and its preimage in equals Suppose that is an open-closed subset such that Then hence for some we have i.e. ∎
@@
Consider the surjective map Put and Then Put Then the diagram is the limit of but is a one-element set for Hence, the sequential limits do not commute with coequalizers in ∎
@@ Appendix G K-theory of products of exact infinity-categories
In this section we prove that -theory of exact -categories commutes with infinite products (Corollary G.9).
@@ G.1. Products of additive infinity-categories
Given a small additive -category we denote by its stabilization. By [Bon, Theorem 5.3.1, Proposition 6.2.1], if is Karoubi complete, then so is Given we denote by the full subcategory generated by We put and similarly The following result is due to A. Córdova Fedeli.
@@
Since we will need an analogous statement for -structures, we recall the proof from [Cor23] with some simplifications. We introduce some notation. Put Given functions we put
@@
[Cor23, Lemma 2.9] 1) For each the functors
have the same class in 2) Analogous statement holds for the same functors
@@
Then and the map becomes an isomorphism in This proves that is essentially surjective. Let now be objects of and denote by and their images in Fullness of Consider a map represented by a roof where and Take such that and Choose and the maps such that Then we can choose factorizations and automatically Putting we have and Therefore, the roof represents a morphism such that Faithfulness of Consider a map such that We may assume that can be lifted to a morphism Then there exists a map such that and Arguing as above, we can construct a factorization of such that and Hence ∎
@@ G.2. Products of stable categories with bounded -structures
We will need the following analogue of Theorem G.1. Let be a family of small stable categories with bounded -structures. Denote by the full subcategory formed by objects such that the objects have uniformly bounded homological amplitude.
@@
The proof is completely analogous to the proof of Theorem G.1 with minor changes, which we indicate. For a stable category with a bounded -structure and for integers we denote by the full subcategory formed by objects whose homology is concentrated in degrees between and and similarly for Put Given functions we define the subcategories analogously to (G.1)-(G.3). Analogously to Lemma G.3, one shows that we have an isomorphism Analogously to Lemma G.4, one proves that for any function the functor is an equivalence. Then one concludes that proving the theorem.∎
@@ G.3. Products of exact -categories
[Bar15, Definition 3.1] An exact -category is a triple where is an additive -category and are subcategories such that the following conditions hold. 1) Every morphism (resp. ) is in (resp. ). 2) Pushouts of morphisms in exist, and the class of morphisms in is closed under pushouts. Pullbacks of morphisms in exist and the class of morphisms in is closed under pullbacks. 3) For a commutative square
the following are equivalent: a) the square is cocartesian, and b) the square is cartesian, and
@@
where and A total cofiber of such a square in is by definition a primitive acyclic object. The full subcategory of acyclic objects is the stable subcategory generated by the primitive acyclic objects. The stable hull of is the category we follow the notation of [Kle22].
@@
Note that if is a triangulated category with a -structure, and is a weak Serre subcategory of the heart, then the full triangulated subcategory generated by has a bounded -structure with the heart In particular, the functor is -exact. Consider the category with its standard -structure. By [Kle22, Proposition 3.10], the primitive acyclic objects are contained in the heart Moreover, by [Kle22, Lemma 3.11 and Lemma 3.12] they form a weak Serre subcategory. The proposition follows.∎
@@
By Theorem G.1, applied to the middle vertical arrow is an isomorphism. By Theorem G.5 applied to the bounded -structures on from Proposition G.7, applied to the left vertical arrow is an isomorphism. Hence, applied to the right vertical arrow is an isomorphism. This proves the theorem.∎